26.003 The refractive indices of materials A and B have a ratio of nA/nB = 1.26....

26.003

The refractive indices of materials A and B have a ratio of nA/nB = 1.26. The speed of light in material A is 1.09 × 108 m/s. What is the speed of light in material B?

26.006

Light has a wavelength of 268.9 nm and a frequency of 7.391 × 1014 Hz when traveling through a certain substance. What is the refraction index of this medium? Use 2.998 × 108 m/s for the speed of light in a vacuum.
Answer #1

26.003

we know, n = c/v

nA/nB = (c/vA)/(c/vB)

nA/nB = vB/vA

==> vB = vA*(nA/nB)

= 1.09*10^8*1.26

= 1.37*10^8 m/s <<<<<<<<<<------------------Answer

26.006
speed of light in the material, v = lamda*f

= 268.9*10^-9*7.391*10^14

= 1.987*10^8 m/s

refractive index of the material, n = c/v

= 2.998*10^8/(1.987*10^8)

= 1.51   <<<<<<<<<<------------------Answer